A solution of 70% alcohol is to be mixed with a solution of 34% alcohol to form 72 liters of a 47% solution. How many liters of the 70% solution must be used?

NEED HELP ASAP!

Answer by ewatrrr(10682)

Hi,
Let x be the total amount of the 70% alcohol solution. then (72L -x) would be the total amount of the 34% alcoohol solution.
.
question states:
.70*x + .34*(72L-x) = .47*72
.
simplify and solve
.70*x + .34*72L- .34*x = .47*72L
.36*x = .47*72L - .34*72L
.36x = 9.36L
x = 26L (the amount of the 70% alcohol solution needed to be used)